Foundations of complex Analysis in non locally Convex Spaces

December 1, 2003

Aboubakr Bayoumi. King Saud University (Saudi Arabia)

Abstract. In fact, real development in Function Theory has been given after we have introduced several new concepts, especially the concept of Quasi-Differentiability which is stronger than Fréchet one. We have built up a complete theory for Function Theory in F-spaces which are not necessarily locally convex.

This lecture will be on a part of our first book entitled: “Foundations of complex Analysis in non locally Convex Spaces” which appear this year.